OpenStudy (anonymous):

find b if v=<1/6,b) is a unit vector

geerky42 (geerky42):

\(\mathbf v = \langle\frac{1}{6},b\rangle\) ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\sqrt{(1/6)^2 + b^2)} =1\] how do i solve for b?

geerky42 (geerky42):

Try squaring both sides.